OpenStudy (anonymous):

Solve on the interval [0, 2π): 2sin^2x+3sinx+1=0 A. x= 3π/2, x=7π/6, x=11π/6 B. x=π,x=2π/3, x=5π/3 C. x=2π, x=π/3 D. x=2π, x=π/4, x= 5π/4

OpenStudy (anonymous):

Lets x = sinx \[2x^2 + 3x +1 = 0\] \[(2x +1)(x +1)=0\] \[x = \frac{-1}{2} , -1\]

OpenStudy (anonymous):

\[sinx = \frac{-1}{2},-1\] when x = 3pi/2 ,sinx =-1 So A

OpenStudy (anonymous):

option A

OpenStudy (anonymous):

did u understand @kassia